What is the volume if a pyramid with a height of 30 m and a sqaure base whose sides are 30 m each?
Answer by rapaljer(4671)

V=(1/3)*B*h, where B= Area of Base = 30^2= 900 m^2, and h= height=30 m
V=(1/3)*900m^2*30 m
V=9000 m^3

Dr. Robert J. Rapalje